OpenStudy (anonymous):

Tamara has decided to start saving for spending money for her first year of college. Her money is currently in a large suit case under her bed, modeled by the function s(x) = 450. She is able to babysit to earn extra money and that function would be a(x) = 6(x − 2), where x is measured in hours. Explain to Tamara how she can create a function that combines the two and describe any simplification that can be done. 450+6(x-2) This is what i got but that would be wrong because there was no simplifying.

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

@primeralph

OpenStudy (anonymous):

if i simplify would that make it different?

OpenStudy (shamil98):

450 + 6(x-2) 450 + 6x -12 438 + 6x

OpenStudy (anonymous):

ok see i was thinking of simplifying it to 450+6x-12 and keeping it that way and I thought it would mess it up thank you.

OpenStudy (shamil98):

actually wait..

OpenStudy (shamil98):

Her money is 450 she earns money for f(x) = 6(x-2) f(x)= 6x - 12

OpenStudy (anonymous):

yah wait that would mess it up because 450 should not be altered due to its in the suitcase

OpenStudy (anonymous):

but you have to combine the total of 450 and the extra money she earns into one equation

OpenStudy (shamil98):

450 + (6x-12) 450 + (12 -12) 450 + 6(2) - 12 is the same I would leave it as (6x-12) + 450 = f(x)

OpenStudy (anonymous):

would it be the same as f(x) = 450+(6x+12)

OpenStudy (shamil98):

no +12 and -12 are different

OpenStudy (shamil98):

if that is what you are asking.. ?..

OpenStudy (anonymous):

nvm